Find the area of the region bounded by y= 11x2/16, y=x2, y=3-2x, satisfying x≤0





.


1
Expert's answer
2022-05-02T16:00:06-0400

Schematically plot region



Area A = A1 + A2


Parts

"A1 = \int_{x1}^{x2}\int_{11/16x^2}^{3-2x}dydx"

"A2 = \int_{x2}^{0}\int_{11/16x^2}^{x^2}dydx"


Integration boundaries

"x1:\\
3 - 2x_1 = 11/16x_1^2, x_1 \le 0\\
x_1 = -4"


"x2:\\

3-2x_2 = x_2^2, x_2 \le 0\\

x_2 = -3"


Area parts

"A1 = \int_{-4}^{-3}\int_{11/16x^2}^{3-2x}dydx\\

= \int_{-4}^{-3} 3 - 2x - 11/16x^2dx\\

= \Big[ 3x - x^2 - 11/48x^3 \Big]_{-4}^{-3}\\

= 3*(-3 + 4) - (9 - 16) - 11/48*(-27 + 64) \approx 1.5208"


"A2 = \int_{-3}^{0}\int_{11/16x^2}^{x^2}dydx\\

= \int_{-3}^{0} x^2 - 11/16x^2dx=\int_{-3}^{0} 5/16x^2dx\\

= 5/48x^3\Big|_{-3}^{0}\\

= 5/48*(0 + 27) = 2.8125"


Area

"A = A1 + A2 \approx1.5208 + 2.8125 = 4.3333"